OpenStudy (anonymous):

a costumer went to a garden a shop and bought some potting soil for $12.50 and 6 shurbs if the total bill was 71.00 how much did the shrubs cost if they all cost the same write and solve an equation show your work

OpenStudy (anonymous):

x=(71-12.50)/6 x=58.5/6 x=9.75 Each shrub cost $9.75

OpenStudy (anonymous):

ok as you can see the total bill is equal to 12,5+6x=71 where x is the number of shurbs, so you obtain 6x=71-12.5=58.5, that is 6x=58.5, dividing by six you have x= 9.75

OpenStudy (thread252):

shrub cost 9.75 dollar

OpenStudy (anonymous):

9.75 each! 71.00-12.50 = 58.50 58.50/6 = 9.75 check: 9.75x6=58.50+12.50 = 71.00